A high school graduating class is made of 365 students. There are 111 more girls than boys. How many boys are in the class? Pleas show me step by step the process of the problem.
Found 2 solutions by rfer, nyc_function:
Answer by rfer(16322)

x+x+111=365
2x=254
x=127 boys
x+111=238 girls

Answer by nyc_function(2741)

Let b = boys
Let g = girls
b + g = 365
111 + g = b
Can you finish?